(10 pts) 1. Evaluate or determine the definite integral: Ni (a) (b) Seat (c) / xe dt. sin(at)dt.
Answer #1

dt tt .4C n+1 | Jach doc=anti = 3 th d t = thel :), ET Cat-6) (6-0 IN 33 3w3 303 17 dt = 1/3 - 2010 je atos-earth 3N3 y et de